OpenStudy (anonymous):

PLEZ HELP ME Joyce combines 9 atoms of hydrogen and 3 atoms of nitrogen to make 3 molecules of ammonia. write a direct variation equation which relates atoms of hydrogen to atoms of nitrogen.

ganeshie8 (ganeshie8):

something like below : `9(hydrogen) + 3(nitrogen) = ammonia`

ganeshie8 (ganeshie8):

`9h + 3n = a`

OpenStudy (anonymous):

(1/9)*h*n wud be correct right?

ganeshie8 (ganeshie8):

`9h = 3n` `h = (3/9)n`

OpenStudy (anonymous):

life saver thanks!!

ganeshie8 (ganeshie8):

yw